# Many-body quantum catalysts for transforming between phases of matter

### v1

@@ Abstract
A catalyst is a substance that enables otherwise impossible transformations between states of a system, without being consumed in the process. In this work, we apply the notion of catalysts to many-body quantum physics. In particular, we construct catalysts that enable transformations between different symmetry-protected topological (SPT) phases of matter using symmetric finite-depth quantum circuits. We discover a wide variety of catalysts, including GHZ-like states which spontaneously break the symmetry, gapless states with critical correlations, topological orders with symmetry fractionalization, and spin-glass states. These catalysts are all united under a single framework which has close connections to the theory of quantum anomalies, and we use this connection to put strong constraints on possible pure- and mixed-state catalysts. We also show how the catalyst approach leads to new insights into the structure of certain phases of matter, and to new methods to efficiently prepare SPT phases with long-range interactions or projective measurements.
@@ I Introduction
One can additionally impose symmetries on the Hamiltonian or circuit, and a state that can be mapped to a product state using an FDQC, but cannot be using a symmetric FDQC, is said to belong to a non-trivial symmetry-protected topological (SPT) phase of matter Chen et al. (2013). This is the class of states that we focus on in this work. In particular, we construct catalysts that enable transformations between distinct SPT phases of matter using symmetric FDQCs, something that is impossible without a catalyst, see Fig. 1. These catalysts have widely varying physical properties and include states like symmetry-breaking states with cat-like entanglement, gapless ground states of conformal field theories, topologically ordered states with fractionalized excitations, and disordered spin-glass-like states. All of these types of catalyst are unified under a single framework. Namely, if a transformation between SPT phases with symmetry group $G$ can be implemented by a symmetric quantum cellular automaton $\mathcal{U}$, then any state that is invariant under both $G$ and $\mathcal{U}$ can catalyze the transformation by symmetric FDQC. Remarkably, the combination of $G$ and $\mathcal{U}$ symmetry has deep connections to the notion of quantum anomalies which, in particular, implies that pure-state catalysts must have long-range entanglement (although we show this is not true for mixed-state catalysts).
Using our framework, we construct explicit examples of catalysts and their associated symmetric FDQCs to transformation between a large variety of SPT phases of matter. In some cases, the form of the catalysts helps us to learn more about the SPT phase at hand. In other cases, the catalyzed transformation can be used to construct new methods of efficient state preparation using, for example, long-range power-law decaying interactions or projective measurements. Overall, our results pave the way for studying quantum catalysts in the context of many-body quantum systems.
@@ II.1 Definitions
$\mathcal{U}=\prod_{\ell=1}^{D}\left(\prod_{i}u_{\ell,i}\right)$ (1)
@@
In this paper, we focus mainly on bosonic SPT phases with an on-site unitary symmetry $U(g)=\prod_{i}u_{i}(g)$ for $g\in G$ (which is not necessarily spatially uniform). The classification of SPT phases is based on equivalence classes of quantum states modulo symmetric FDQCs Chen et al. (2013). That is, two states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ are in the same SPT phase if and only if there is a symmetric FDQC $\mathcal{U}$ such that
@@
$\mathcal{V}(|\psi_{0}\rangle\otimes|a\rangle)=|\psi_{1}\rangle\otimes|a\rangle,$ (3)
Observe that we only consider catalysts which are themselves symmetric such that the symmetry of the whole system is preserved. Here we considered only pure-state catalysts, but in Sec. V will also allow mixed-state catalysts which are defined using a suitable generalization of Definition 2. Naturally, since the definition of many-body catalysts include the notion of an FDQC, it should only be applied in a scenario where the initial, final, and catalyst states can be defined on systems of increasing size, and the depth $D$ and range $r$ of the FDQC are both independent of system size.
@@
This operational definition also leads to a strong constraint on the form of possible catalysts. Namely, it cannot be possible to generate a catalyst for a non-trivial SPT phase by acting with a symmetric FDQC on a product state. If this was possible, then the process of creating the catalyst, using it to catalyze the SPT phase, and destroying the catalyst describes an symmetric FDQC that creates the SPT phase from a product state, which is a contradiction. More generally, the catalyst $|a\rangle$ cannot be symmetrically invertible Lessa et al. (2024a), meaning that there cannot exist a state $|a^{-1}\rangle$ such that $|a\rangle\otimes|a^{-1}\rangle$ is connected to a symmetric product state via a symmetric FDQC. In other words, it must at least as hard to make the catalyst as it is to make the SPT phase on its own. However, the advantage of using a catalyst is that the hard work only needs to be performed once, since the catalyst is not consumed after use and can therefore be re-used.
@@ II.2 $(G,\mathcal{U})$-symmetric catalysts
$\mathcal{V}(|\psi\rangle\otimes|a\rangle)=\mathcal{U}|\psi\rangle\otimes\mathcal{U}^{-1}|a\rangle=\mathcal{U}|\psi\rangle\otimes|a\rangle$ (6)
@@
Theorem 1 is an important result that will help us to construct a large class of SPT-catalysts with various physical properties. Furthermore, Theorem 1 establishes an important relationship between catalysts and anomalies. In the present context, a symmetry group is said to be anomalous if no short-range entangled state can be invariant under all elements of the group. A state is said to have short-range entanglement (SRE) if it can be prepared from a product state using an FDQC, and otherwise it has long-range entanglement (LRE) Chen et al. (2010). In general, we expect that the set of symmetries $\mathcal{U}$ and $U(g)$ for all $g\in G$ is anomalous whenever $\mathcal{U}$ is an entangler for a non-trivial SPT phase. One can argue this based on the fact that these are the symmetries that emerge on the boundary of a decorated domain wall SPT-ordered state in one higher dimension Chen et al. (2014), and it is known that such states do not allow short-range entangled boundary states. For the case of abelian symmetries in 1D systems, we explicitly prove in Sec. V that $(G,\mathcal{U})$-symmetric catalysts must have LRE.
@@ II.3 Case study: 1D SPT and LSM
$|\mathrm{triv.}\rangle=\bigotimes_{i=1}^{N}|\Omega\rangle_{2i,2i+1}$ (7)
$|\mathrm{SPT}\rangle=\bigotimes_{i=1}^{N}|\Omega\rangle_{2i-1,2i}$ (8)
@@
$H_{a}=H_{\mathrm{triv.}}+H_{\mathrm{SPT}}$ (13)
@@
This example is interesting because it leads to a novel method to prepare SPT phases with long-range interactions, as discussed in Sec. VI.
@@ III Catalysts for SPT phases with global symmetries
$|\psi(\nu_{d})\rangle=\mathcal{U}(\nu_{d})|+_{G}+_{G}\cdots+_{G}\rangle$ (15)
@@
$|a\rangle=\frac{1}{\mathcal{N}}\sum_{\ell=1}^{L}\mathcal{U}(\nu_{d})^{\ell}|+_{G}+_{G}\cdots+_{G}\rangle.$ (17)
@@
$H_{a}=\sum_{\ell=1}^{L}\mathcal{U}(\nu_{d})^{\ell}H_{\mathrm{triv.}}\mathcal{U}(\nu_{d})^{\ell\dagger}$ (18)
@@
$H_{a}(\alpha)=\alpha H_{\mathrm{triv.}}+(1-\alpha)H_{\mathrm{SPT}}$ (19)
@@ III.1 Example: 1D cluster state
$|\psi_{C}\rangle=\mathcal{U}_{\mathrm{CZ}}|++\cdots+\rangle$ (20)
@@
$\mathcal{U}_{\mathrm{CZ}}=\prod_{i=1}^{N}{\mathrm{CZ}}_{i,i+1}$ (21)
@@
$H(\alpha)=-\alpha\sum_{i}X_{i}-(1-\alpha)\sum_{i}Z_{i-1}X_{i}Z_{i+1}$ (22)
@@ III.2 Topologically ordered catalysts in $d>1$
$|\mathrm{SPT}\rangle=\mathcal{U}_{\mathrm{SPT}}\left(|0\rangle^{\otimes|E|}\otimes|+\rangle^{\otimes|V|}\right)$ (25)
@@
$\mathcal{U}_{\mathrm{SPT}}=\prod_{v}\mathrm{ctrl}-A_{v}$ (26)
@@
$\mathcal{U}_{\mathrm{SPT}}|\widetilde{DS}\rangle=\sum_{\vec{i}}\left(\prod_{v}A_{v}^{i_{v}}\right)|DS\rangle\otimes|\vec{i}\rangle=|\widetilde{DS}\rangle$ (28)
@@ IV.1 Higher-form symmetries
$|\psi_{\mathrm{Lieb}}\rangle=\mathcal{U}_{\mathrm{Lieb}}\left(|+\rangle^{\otimes|E|}\otimes|+\rangle^{\otimes|V|}\right)$ (30)
@@
$\mathcal{U}_{\mathrm{Lieb}}=\prod_{\langle e,v\rangle}CZ_{e,v}$ (31)
@@
$|a\rangle=|+\rangle^{\otimes|E|}\otimes\left(|0\rangle^{\otimes|V|}+|1\rangle^{\otimes|V|}\right)$ (33)
@@
$|a\rangle=\left(\prod_{p}\frac{1+B_{p}}{2}|0\rangle^{\otimes|E|}\right)\otimes|+\rangle^{\otimes|V|}$ (34)
@@ IV.2 Subsystem symmetries
$|\psi_{\mathrm{square}}\rangle=\mathcal{U}_{\mathrm{square}}|+\rangle^{|V|}$ (36)
@@
$H_{PIM}=-\sum_{v}\left(\prod_{v^{\prime}\in n(v)}Z_{v^{\prime}}\right)$ (39)
@@ V Mixed state catalysts
This means that the state $\rho$ is a mixture of symmetric states, rather than only being symmetric “on average”, such that the symmetry of the system is always preserved. If we allow weakly-symmetric states which don’t satisfy Eq. (40) but do satisfy $U(g)\rho U(g)^{\dagger}=\rho$, then it is trivial to construct catalysts, the maximally mixed state being one example. We then define a mixed state catalyst in a similar fashion as for pure states,
@@
$\mathcal{V}(|\psi_{0}\rangle\langle\psi_{0}|\otimes\rho)\mathcal{V}^{\dagger}=|\psi_{1}\rangle\langle\psi_{1}|\otimes\rho$ (41)
@@ V.1 Catalysts from strong-to-weak symmetry breaking
$\begin{gathered}\mathrm{Tr}\left(\rho W_{\lambda}\right)=0\\ F\left(\rho,W_{\lambda}\rho W_{\lambda}\right)=1\end{gathered}.$ (50)
@@
$\begin{gathered}\mathrm{Tr}\left(\rho U^{(1)}_{\bar{\gamma}}\right)=0\\ F\left(\rho,U^{(1)}_{\bar{\gamma}}\rho U^{(1)}_{\bar{\gamma}}\right)=1\end{gathered}.$ (53)
@@ V.2 Characterizing 1D catalysts
$\rho U_{\Gamma,g}=\rho(L_{\Gamma,g}\otimes R_{\Gamma,g})$ (54)
@@
$\mathcal{U}^{\dagger}U_{A,g}\mathcal{U}U_{B,h}\mathcal{U}^{\dagger}U_{A,g}^{\dagger}\mathcal{U}U_{B,h}^{\dagger}=c_{g,h}\mathbb{1}$ (55)
@@
$\tilde{\mathcal{W}}_{\Gamma,g}$ $=\mathcal{U}^{\dagger}W_{\Gamma,g}\mathcal{U}$ (58)
@@
$=c_{g,h}\mathrm{Tr}(\mathcal{U}^{\dagger}U_{A,g}\mathcal{U}U_{B,h}\mathcal{U}^{\dagger}\mathcal{W}_{B,h}^{\dagger}U_{A,g}^{\dagger}\rho\mathcal{U})$
@@
$=c_{g,h}c_{h^{-1},g}\mathrm{Tr}(\mathcal{U}^{\dagger}U_{A,g}\mathcal{U}\tilde{\mathcal{W}}_{A,g}^{\dagger}\rho)$
@@
$U_{\Gamma,g}^{\dagger}\rho U_{\Gamma,g}=(L_{\Gamma,g}^{\dagger}\otimes R_{\Gamma,g}^{\dagger})\rho(L_{\Gamma,g}\otimes R_{\Gamma,g})$ (62)
@@
$\mathcal{U}U_{\Gamma,g}(L_{\Gamma,g}^{\dagger}\otimes R_{\Gamma,g}^{\dagger})\mathcal{U}^{\dagger}=U_{\Gamma,g}(\mathcal{L}_{\Gamma,g}^{\dagger}\otimes\mathcal{R}_{\Gamma,g}^{\dagger})$ (63)
@@
$=(\mathcal{L}_{\Gamma,g}L_{\Gamma,g}^{\dagger}\otimes\mathcal{R}_{\Gamma,g}R_{\Gamma,g}^{\dagger})\rho(L_{\Gamma,g}\mathcal{L}_{\Gamma,g}^{\dagger}\otimes R_{\Gamma,g}\mathcal{R}_{\Gamma,g}^{\dagger})$

### v2

@@ Abstract
A catalyst is a substance that enables otherwise impossible transformations between states of a system, without being consumed in the process. In this work, we apply the notion of catalysts to many-body quantum physics. In particular, we construct catalysts that enable transformations between different symmetry-protected topological (SPT) phases of matter using symmetric finite-depth quantum circuits. We discover a wide variety of catalysts, including GHZ-like states that spontaneously break the symmetry, gapless states with critical correlations, topological orders with symmetry fractionalization, and spin-glass states. These catalysts are all united under a single framework that has close connections to the theory of quantum anomalies, and we use this connection to put strong constraints on possible pure- and mixed-state catalysts. We also show how the catalyst approach leads to new insights into the structure of certain phases of matter, and to new methods to efficiently prepare SPT phases with long-range interactions or projective measurements.
@@ I Introduction
One can additionally impose symmetries on the Hamiltonian or circuit, and a state that can be mapped to a product state using an FDQC, but cannot be using a symmetric FDQC, is said to belong to a non-trivial symmetry-protected topological (SPT) phase of matter Chen et al. (2013). This is the class of states that we focus on in this work. In particular, we construct catalysts that enable transformations between distinct SPT phases of matter using symmetric FDQCs, something that is impossible without a catalyst, see Fig. 1. These catalysts are themselves many-body states matching the dimensionality of the system, and they have widely varying physical properties including symmetry-breaking states with cat-like entanglement, gapless ground states of conformal field theories, topologically ordered states with fractionalized excitations, and disordered spin-glass states. All of these types of catalyst are unified under a single framework. Namely, if a transformation between SPT phases with symmetry group $G$ can be implemented by a symmetric quantum cellular automaton $\mathcal{U}$, then any state that is invariant under both $G$ and $\mathcal{U}$ can catalyze the transformation by symmetric FDQC. Remarkably, the combination of $G$ and $\mathcal{U}$ symmetry has deep connections to the notion of quantum anomalies which, in particular, implies that pure-state catalysts must have long-range entanglement (although we show this is not true for mixed-state catalysts).
Using our framework, we construct explicit examples of catalysts and their associated symmetric FDQCs to transform between a large variety of SPT phases of matter with on-site unitary symmetry groups. In general, pure state catalysts are at least as difficult to prepare as the SPT phases themselves. However, since the catalyst can be reused, this hard work only needs to be done once, since a single catalyst can be used to prepare many SPT states. On the other hand, we show that mixed-state catalysts can be prepared in finite time using symmetric local channels. Furthermore in some cases, the form of the catalysts helps us to learn more about physical properties the SPT phase at hand. We also show how the catalyzed transformation can be used to construct new methods of efficient state preparation using, for example, long-range power-law decaying interactions or projective measurements. Overall, our results pave the way for studying quantum catalysts in the context of many-body quantum systems.
@@ II.1 Definitions
$\mathcal{U}=\prod_{\ell=1}^{D}\left(\prod_{i}u_{\ell,i}\right),$ (1)
@@
In this paper, we consider bosonic SPT phases with an on-site unitary symmetry $U(g)=\prod_{i}u_{i}(g)$ for $g\in G$ (which is not necessarily spatially uniform). The objects of interest are symmetric quantum states which satisfy $U(g)|\psi\rangle=|\psi\rangle$ for all $g\in G$. The classification of SPT phases is based on equivalence classes of symmetric quantum states modulo symmetric FDQCs Chen et al. (2013). That is, two symmetric states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ are in the same SPT phase if and only if there is a symmetric FDQC $\mathcal{U}$ such that
@@
$\mathcal{V}(|\psi_{0}\rangle\otimes|a\rangle)=|\psi_{1}\rangle\otimes|a\rangle.$ (3)
Observe that we only consider catalysts which are themselves symmetric such that the symmetry of the whole system is preserved. Furthermore, the dimensionality of the catalyst should match that of the system that is being catalyzed, i.e. the catalyst should be defined on a constant number of ancillary degrees of freedom per site $i$. If one allows higher dimensional catalysts, for example, then the pumping construction of Ref. Tantivasadakarn and Vishwanath (2022) shows that even a symmetric product state can serve as a catalyst.
Here we considered only pure-state catalysts, but in Sec. V will also allow mixed-state catalysts which are defined using a suitable generalization of Definition 2. Naturally, since the definition of many-body catalysts include the notion of an FDQC, it should only be applied in a scenario where the initial, final, and catalyst states can be defined on systems of increasing size, and the depth $D$ and range $r$ of the FDQC are both independent of system size.
@@
This operational definition also leads to a strong constraint on the form of possible catalysts. Namely, it cannot be possible to generate a catalyst for a non-trivial SPT phase by acting with a symmetric FDQC on a product state. If this was possible, then the process of creating the catalyst, using it to catalyze the SPT phase, and destroying the catalyst describes an symmetric FDQC that creates the SPT phase from a product state, which is a contradiction. More generally, the catalyst $|a\rangle$ cannot be symmetrically invertible Lessa et al. (2024a), meaning that there cannot exist a state $|a^{-1}\rangle$ such that $|a\rangle\otimes|a^{-1}\rangle$ is connected to a symmetric product state via a symmetric FDQC. In other words, it must be at least as hard to make the catalyst as it is to make the SPT phase on its own. However, the advantage of using a catalyst is that the hard work only needs to be performed once, since the catalyst is not consumed after use and can therefore be re-used.
@@ II.2 $(G,\mathcal{U})$-symmetric catalysts
$\mathcal{V}(|\psi\rangle\otimes|a\rangle)=\mathcal{U}|\psi\rangle\otimes\mathcal{U}^{-1}|a\rangle=\mathcal{U}|\psi\rangle\otimes|a\rangle,$ (6)
@@
Theorem 1 is an important result that will help us to construct a large class of SPT-catalysts with various physical properties. Furthermore, Theorem 1 establishes an important relationship between catalysts and anomalies. In the present context, a symmetry group is said to be anomalous if no short-range entangled state can be invariant under all elements of the group. A state is said to have short-range entanglement (SRE) if it can be prepared from a product state using an FDQC, and otherwise it has long-range entanglement (LRE) Chen et al. (2010). In general, we expect that the set of symmetries $\mathcal{U}$ and $U(g)$ for all $g\in G$ is anomalous whenever $\mathcal{U}$ is an entangler for a non-trivial SPT phase. One can argue this based on the fact that these are the symmetries that emerge on the boundary of a decorated domain wall SPT-ordered state in one higher dimension Chen et al. (2014), and it is known that such states do not allow short-range entangled boundary states. For the case of abelian symmetries in 1D systems, we explicitly prove in Sec. V that $(G,\mathcal{U})$-symmetric catalysts must have LRE. The fact that anomalies can “eat” SPT phases has previously been observed in Refs. Hsin and Turzillo (2020); Dumitrescu and Hsin (2024); Choi et al. (2024).
@@ II.3 Case study: 1D SPT and LSM
$|\mathrm{triv.}\rangle=\bigotimes_{i=1}^{N}|\Omega\rangle_{2i,2i+1},$ (7)
$|\mathrm{SPT}\rangle=\bigotimes_{i=1}^{N}|\Omega\rangle_{2i-1,2i},$ (8)
@@
$H_{a}=H_{\mathrm{triv.}}+H_{\mathrm{SPT}},$ (13)
@@
This example is interesting because it leads to a novel method to prepare SPT phases with long-range interactions, as discussed in Sec. VI. Although this example does not fit into the LSM framework, it is constrained by a more general condition that we describe in Sec. V (Theorem 2).
@@ III Catalysts for SPT phases with global symmetries
$|\psi(\nu_{d})\rangle=\mathcal{U}(\nu_{d})|+_{G}+_{G}\cdots+_{G}\rangle,$ (15)
@@
$|a\rangle=\frac{1}{\mathcal{N}}\sum_{\ell=1}^{L}\mathcal{U}(\nu_{d})^{\ell}|+_{G}+_{G}\cdots+_{G}\rangle,$ (17)
@@
$H_{a}=\sum_{\ell=1}^{L}\mathcal{U}(\nu_{d})^{\ell}H_{\mathrm{triv.}}\mathcal{U}(\nu_{d})^{\ell\dagger}.$ (18)
@@
$H_{a}(\alpha)=\alpha H_{\mathrm{triv.}}+(1-\alpha)H_{\mathrm{SPT}},$ (19)
@@ III.1 Example: 1D cluster state
$|\psi_{C}\rangle=\mathcal{U}_{\mathrm{CZ}}|++\cdots+\rangle,$ (20)
@@
$\mathcal{U}_{\mathrm{CZ}}=\prod_{i=1}^{N}{\mathrm{CZ}}_{i,i+1},$ (21)
@@
$H(\alpha)=-\alpha\sum_{i}X_{i}-(1-\alpha)\sum_{i}Z_{i-1}X_{i}Z_{i+1},$ (22)
@@ III.2 Topologically ordered catalysts in $d>1$
$|\mathrm{SPT}\rangle=\mathcal{U}_{\mathrm{SPT}}\left(|0\rangle^{\otimes|E|}\otimes|+\rangle^{\otimes|V|}\right),$ (25)
@@
$\mathcal{U}_{\mathrm{SPT}}=\prod_{v}\mathrm{ctrl}-A_{v},$ (26)
@@
$\mathcal{U}_{\mathrm{SPT}}|\widetilde{DS}\rangle=\sum_{\vec{i}}\left(\prod_{v}A_{v}^{i_{v}}\right)|DS\rangle\otimes|\vec{i}\rangle=|\widetilde{DS}\rangle,$ (28)
@@ IV.1 Higher-form symmetries
$|\psi_{\mathrm{Lieb}}\rangle=\mathcal{U}_{\mathrm{Lieb}}\left(|+\rangle^{\otimes|E|}\otimes|+\rangle^{\otimes|V|}\right),$ (30)
@@
$\mathcal{U}_{\mathrm{Lieb}}=\prod_{\langle e,v\rangle}CZ_{e,v}.$ (31)
@@
$|a\rangle=|+\rangle^{\otimes|E|}\otimes\left(|0\rangle^{\otimes|V|}+|1\rangle^{\otimes|V|}\right),$ (33)
@@
$|a\rangle=\left(\prod_{p}\frac{1+B_{p}}{2}|0\rangle^{\otimes|E|}\right)\otimes|+\rangle^{\otimes|V|},$ (34)
@@ IV.2 Subsystem symmetries
$|\psi_{\mathrm{square}}\rangle=\mathcal{U}_{\mathrm{square}}|+\rangle^{|V|},$ (36)
@@
$H_{PIM}=-\sum_{v}\left(\prod_{v^{\prime}\in n(v)}Z_{v^{\prime}}\right).$ (39)
@@ V Mixed state catalysts
This means that the state $\rho$ is a mixture of symmetric states, rather than only being symmetric “on average”, such that the symmetry of the system is always preserved. If we allow weakly-symmetric states which don’t satisfy Eq. (40) but do satisfy $U(g)\rho U(g)^{\dagger}=\rho$, then it is trivial to construct catalysts, the maximally mixed state being one example. This matches the known fact that SPT order is not stable under weakly-symmetric channels de Groot et al. (2022). We then define a mixed state catalyst in a similar fashion as for pure states,
@@
$\mathcal{V}(|\psi_{0}\rangle\langle\psi_{0}|\otimes\rho)\mathcal{V}^{\dagger}=|\psi_{1}\rangle\langle\psi_{1}|\otimes\rho.$ (41)
@@ V.1 Catalysts from strong-to-weak symmetry breaking
$\begin{gathered}\mathrm{Tr}\left(\rho W_{\lambda}\right)=0\\ F\left(\rho,W_{\lambda}\rho W_{\lambda}\right)=1.\end{gathered}$ (50)
@@
$\begin{gathered}\mathrm{Tr}\left(\rho U^{(1)}_{\bar{\gamma}}\right)=0\\ F\left(\rho,U^{(1)}_{\bar{\gamma}}\rho U^{(1)}_{\bar{\gamma}}\right)=1.\end{gathered}$ (53)
@@ V.2 Characterizing 1D catalysts
$\rho U_{\Gamma,g}=\rho(L_{\Gamma,g}\otimes R_{\Gamma,g}).$ (54)
@@
$\mathcal{U}^{\dagger}U_{A,g}\mathcal{U}U_{B,h}\mathcal{U}^{\dagger}U_{A,g}^{\dagger}\mathcal{U}U_{B,h}^{\dagger}=c_{g,h}\mathbb{1},$ (55)
@@
$\tilde{\mathcal{W}}_{\Gamma,g}$ $=\mathcal{U}^{\dagger}W_{\Gamma,g}\mathcal{U}.$ (58)
@@
$=c_{g,h}\mathrm{Tr}(\mathcal{U}^{\dagger}U_{A,g}\mathcal{U}U_{B,h}\mathcal{U}^{\dagger}\mathcal{W}_{B,h}^{\dagger}U_{A,g}^{\dagger}\rho\mathcal{U}),$
@@
$=c_{g,h}c_{h^{-1},g}\mathrm{Tr}(\mathcal{U}^{\dagger}U_{A,g}\mathcal{U}\tilde{\mathcal{W}}_{A,g}^{\dagger}\rho).$
@@
$U_{\Gamma,g}^{\dagger}\rho U_{\Gamma,g}=(L_{\Gamma,g}^{\dagger}\otimes R_{\Gamma,g}^{\dagger})\rho(L_{\Gamma,g}\otimes R_{\Gamma,g}).$ (62)
@@
$\mathcal{U}U_{\Gamma,g}(L_{\Gamma,g}^{\dagger}\otimes R_{\Gamma,g}^{\dagger})\mathcal{U}^{\dagger}=U_{\Gamma,g}(\mathcal{L}_{\Gamma,g}^{\dagger}\otimes\mathcal{R}_{\Gamma,g}^{\dagger}),$ (63)
@@
$=(\mathcal{L}_{\Gamma,g}L_{\Gamma,g}^{\dagger}\otimes\mathcal{R}_{\Gamma,g}R_{\Gamma,g}^{\dagger})\rho(L_{\Gamma,g}\mathcal{L}_{\Gamma,g}^{\dagger}\otimes R_{\Gamma,g}\mathcal{R}_{\Gamma,g}^{\dagger}),$